Please help me I am unable to slove this expression.
-9-sqrt108 /3

Found 2 solutions by checkley71, edjones:
Answer by checkley71(8403)

-9-SQRT(108/3)
-9-SQRT36
-9-6=-15 ANSWER.

Answer by edjones(8007)

(-9-sqrt(108))/3=(-9-sqrt(36*3))/3=(-9-6sqrt(3))/3=-3-2sqrt(3)
Ed